# Menger and consonant sets in the Sacks model

### v2

@@ 4. Totally imperfect Menger sets and Sacks forcing
For the proof of Theorem 4.1, we need the following auxiliary result, whose proof is rather standard (see, e.g., [7, Lemma 5.10] for a similar argument) and is left to the reader.
@@
Take $p$ and $F_{n},k_{n},l_{n},y_{\sigma_{n}}$ from Lemma 3.7, applied to $p_{0}$ and $\tau$. Note also that Lemma 3.9(3) ensures that $S=\mathrm{supp}(p)$ and the sequences ${\langle\,k_{n}:n\in\omega\,\rangle},{\langle\,F_{n}:n\in\omega\,\rangle},{\langle\,\Sigma_{n}:n\in\omega\,\rangle}$ satisfy $(e_{f})$ from the first paragraph of Section 2. Let $K$ and $[\sigma]$ for $\sigma\in\bigcup_{n\in\omega}\Sigma_{n}$ be defined in the same way as in the first paragraph of Section 2.
@@
$p|\sigma_{n+1}\Vdash\tau\restriction l_{n}=y_{\sigma_{n}}\text{ and }p|\sigma_{n+1}\Vdash\tau\restriction l_{n+1}=y_{\sigma_{n+1}},$
@@ 5. A modification of the Menger game and consonant spaces
Fix a natural number $n>0$ and assume that natural numbers $l_{0},\dotsc,l_{n-1}>0$ and $n-1$ rounds of the game have been defined.
@@
Let $\sigma$ be a strategy for Alice in the grouped Menger game played on $X:={2^{\omega}}\setminus Y$. By Remark 5.1, we may assume that each family given by Alice according to the strategy $\sigma$ is countable and increasing and Bob chooses one set from the families given by Alice. We shall define a strategy §$\S$ for Alice in the game $\mathsf{G}_{1}(\mathcal{K},\mathcal{O})$ played on $X$ such that each play in $\mathsf{G}_{1}(\mathcal{K},\mathcal{O})$ played according to §$\S$ and lost by Alice, gives rise to a play in the grouped Menger game on $X$ in which Alice uses $\sigma$ and loses.
Suppose that $\sigma$ instructs Alice to start round 00 by selecting a natural number $l_{0}>0$. Let $L_{0}:=0$ and $L_{1}:=l_{0}$. Then §$\S$ instructs Alice to play the family of all sets
@@
is a play, where Alice uses the strategy $\sigma$. Note that the family of all the intersections as above is indeed an open $k$-cover of $X$. Suppose that Bob replies in $\mathsf{G}_{1}(\mathcal{K},\mathcal{O})$ by selecting $\bigcap_{i\in[L_{0},L_{1})}U_{i}$ for some sequence $(\mathcal{U}_{0},U_{0},\dotsc,\mathcal{U}_{l_{0}-1},U_{l_{0}-1})$ as above. The strategy $\sigma$ instructs Alice to proceed in round $1$ by selecting a natural number $l_{1}>0$. Let $L_{2}:=l_{0}+l_{1}$. Then §$\S$ instructs Alice to play the family of all sets
@@
In general, let $\sigma$ instruct Alice to start round $n$ by selecting a natural number $l_{n}>0$. Let $L_{n+1}:=l_{0}+l_{1}+\dotsb+l_{n}$. Then the next move of Alice in $\mathsf{G}_{1}(\mathcal{K},\mathcal{O})$ according to §$\S$ is, by the definition, the family of all sets
@@
Since the strategy §$\S$ is not winning, there is a play in $\mathsf{G}_{1}(\mathcal{K},\mathcal{O})$ in which Alice uses §$\S$ and loses, which gives rise to an infinite play
@@ 5.1. Menger game versus the grouped Menger game
Combining [11, Theorem 1.1] and [8, Theorem 10], we conclude that $\mathfrak{d}=\mathfrak{c}$ implies the existence of a Menger ultrafilter.
@@
Let us note that there are models of ZFC without Menger ultrafilters. Indeed, every Menger ultrafilter is a $P$-point, see, e.g., [6, Observation 3.4]. According to a result of Shelah published in [31] (see also [9]), consistently there are no $P$-points.
Theorems 4.1 and 6.1 motivate the following problem.
Problem 5.6.
In the Sacks model:
  1. (1)
  2. (2)
  3. (3)
    Are there only $\omega_{2}$-many Menger subsets of $2^{\omega}$?
  4. (4)
Regarding the last item of Problem 5.6, we do not know the answer even to the following question.
Problem 5.7.
Next, we introduce a variation of the grouped Menger game which is “harder” for Alice since her choices are more restricted, i.e., are not arbitrary open covers of the space in question. Let $K\subseteq{2^{\omega}}$ be a perfect set. For $y\in K$, let $\mathcal{U}_{y}$ be a countable increasing family of clopen sets in $K$ such that $\bigcup\mathcal{U}_{y}=K\setminus\{y\}$. Let $X\subseteq K$ be a set such that $K\setminus X$ is dense in $K$. The weak grouped Menger game played on $X$ in $K$ ($\operatorname{wgM}(K,X)$ in short) is played as follows: In round 00 Alice selects a natural number $l_{0}>0$, and Bob selects a closed nowhere dense subset $K_{0}$ of $K$. Let $L_{0}:=0$ and $L_{1}:=l_{0}$. Then the players play the usual Menger game $l_{0}$ subrounds, with the following restrictions. In each subround $i\in[L_{0},L_{1})$, Alice chooses $y_{i}\in K\setminus(X\cup K_{0})$ and plays the family $\mathcal{U}_{i}:=\mathcal{U}_{y_{i}}$. Then Bob replies by choosing a set $U_{i}\in\mathcal{U}_{i}$ with $K_{0}\subseteq U_{i}$.
@@
[…]
@@
We shall describe a strategy §$\S$ for Alice in the weak grouped Menger game played on $X$ in $K$ such that each play lost Alice gives rise to the objects whose existence we need to establish.
Round 0. Let $C_{0}:=\Sigma_{0}$ and $i_{0}:=0$. Alice declares that the 00-th group will have length $l_{0}:=|C_{0}|$. Let $\{\nu_{j}:j<l_{0}\}\subseteq\Sigma_{0}$ be an enumeration of $C_{0}$. Suppose that Bob plays a closed nowhere dense $K_{0}\subseteq K$.
Subround $\langle 0,0\rangle$. By density of $K\setminus X$, Alice picks $y_{\langle 0,0\rangle}\in[\nu_{0}]\setminus(X\cup K_{0})$ and she plays the family $\mathcal{U}_{\langle 0,0\rangle}:=\mathcal{U}_{y_{\langle 0,0\rangle}}$. Suppose that Bob replies by choosing $U_{\langle 0,0\rangle}\in\mathcal{U}_{\langle 0,0\rangle}$ with $K_{0}\subseteq U_{\langle 0,0\rangle}$. Take $j_{\langle 0,0\rangle}>i_{0}$ and $\nu_{\langle 0,0\rangle}\in\Sigma_{j_{\langle 0,0\rangle}}$ with $\nu_{\langle 0,0\rangle}\succ\nu_{0}$ such that $y_{\langle 0,0\rangle}\in[\nu_{\langle 0,0\rangle}]$ and $[\nu_{\langle 0,0\rangle}]\cap U_{\langle 0,0\rangle}=\emptyset$. Let $e_{\langle 0,0\rangle}\colon C_{0}\to\Sigma_{j_{\langle 0,0\rangle}}$ be a coherent map such that $e_{\langle 0,0\rangle}(\nu_{0})=\nu_{\langle 0,0\rangle}$. Fix a natural number $a$ with $0<a<l_{0}$ and assume that the players have already defined the following sequences:
@@
  • ${\langle\,\nu_{\langle 0,b\rangle}:0\leq b<a\,\rangle}$ of maps $\nu_{\langle 0,b\rangle}\in\Sigma_{j_{\langle 0,b\rangle}}$ such that $y_{\langle 0,b\rangle}\in[\nu_{\langle 0,b\rangle}]$, $\nu_{b}\prec\nu_{\langle 0,b\rangle}$ and $[\nu_{\langle 0,b\rangle}]\cap U_{\langle 0,b\rangle}=\emptyset$,
@@
Subround $\langle 0,a\rangle$: Alice picks $y_{\langle 0,a\rangle}\in[e_{\langle 0,a-1\rangle}(\nu_{a})]\setminus(X\cup K_{0})$ and she plays the family $\mathcal{U}_{\langle 0,a\rangle}:=\mathcal{U}_{y_{\langle 0,a\rangle}}$. Suppose that Bob replies by choosing $U_{\langle 0,a\rangle}\in\mathcal{U}_{\langle 0,a\rangle}$ with $K_{0}\subseteq U_{\langle 0,a\rangle}$. Take $j_{\langle 0,a\rangle}>j_{\langle 0,a-1\rangle}$ and $\nu_{\langle 0,a\rangle}\in\Sigma_{j_{\langle 0,a\rangle}}$ with $\nu_{\langle 0,a\rangle}\succ e_{\langle 0,a-1\rangle}(\nu_{a})$ such that $y_{\langle 0,a\rangle}\in[\nu_{\langle 0,a\rangle}]$ and $[\nu_{\langle 0,a\rangle}]\cap U_{\langle 0,a\rangle}=\emptyset$. Let $e^{\prime}_{\langle 0,a\rangle}\colon e_{\langle 0,a-1\rangle}[C_{0}]\to\Sigma_{j_{\langle 0,a\rangle}}$ be a coherent map such that $e^{\prime}_{\langle 0,a\rangle}(e_{\langle 0,a-1\rangle}(\nu_{a}))=\nu_{\langle 0,a\rangle}$. Put $e_{\langle 0,a\rangle}:=e^{\prime}_{\langle 0,a\rangle}\circ e_{\langle 0,a-1\rangle}$.
@@
satisfy all relevant instances of $(1)$$(4)$.
@@
Subround $\langle n,0\rangle$: Alice picks $y_{\langle n,0\rangle}\in[\nu_{0}]\setminus(X\cup K_{n})$ and she plays the family $\mathcal{U}_{\langle n,0\rangle}:=\mathcal{U}_{y_{\langle n,0\rangle}}$. Suppose that Bob replies by choosing $U_{\langle n,0\rangle}\in\mathcal{U}_{\langle n,0\rangle}$ with $K_{n}\subseteq U_{\langle n,0\rangle}$. Take $j_{\langle n,0\rangle}>i_{n}$ and $\nu_{\langle n,0\rangle}\in\Sigma_{j_{\langle n,0\rangle}}$ with $\nu_{\langle n,0\rangle}\succ\nu_{0}$ such that $y_{\langle n,0\rangle}\in[\nu_{\langle n,0\rangle}]$ and $[\nu_{\langle n,0\rangle}]\cap U_{\langle n,0\rangle}=\emptyset$. Let $e_{\langle n,0\rangle}\colon C_{n}\to\Sigma_{j_{\langle n,0\rangle}}$ be a coherent map such that $e_{\langle n,0\rangle}(\nu_{0})=\nu_{\langle n,0\rangle}$.
@@
  • ${\langle\,\nu_{\langle n,b\rangle}:0\leq b<a\,\rangle}$ of maps $\nu_{\langle n,b\rangle}\in\Sigma_{j_{\langle n,b\rangle}}$ such that $y_{\langle n,b\rangle}\in[\nu_{\langle n,b\rangle}]$, $\nu_{b}\prec\nu_{\langle n,b\rangle}$ and $[\nu_{\langle n,b\rangle}]\cap U_{\langle n,b\rangle}=\emptyset$;
@@
Subround $\langle n,a\rangle$: Alice picks $y_{\langle n,a\rangle}\in[e_{\langle n,a-1\rangle}(\nu_{a})]\setminus(X\cup K_{n})$ and she plays the family $\mathcal{U}_{\langle n,a\rangle}:=\mathcal{U}_{y_{\langle n,a\rangle}}$. Suppose that Bob replies by choosing $U_{\langle n,a\rangle}\in\mathcal{U}_{\langle n,a\rangle}$ such that $K_{n}\subseteq U_{\langle n,a\rangle}$. Take $j_{\langle n,a\rangle}>j_{\langle n,a-1\rangle}$ and $\nu_{\langle n,a\rangle}\in\Sigma_{j_{\langle n,a\rangle}}$ with $\nu_{\langle n,a\rangle}\succ e_{\langle n,a-1\rangle}(\nu_{a})$ such that $y_{\langle n,a\rangle}\in[\nu_{\langle n,a\rangle}]$ and $[\nu_{\langle n,a\rangle}]\cap U_{\langle n,a\rangle}=\emptyset$. Let $e^{\prime}_{\langle n,a\rangle}\colon e_{\langle n,a-1\rangle}[C_{n}]\to\Sigma_{j_{\langle n,a\rangle}}$ be a coherent map such that $e^{\prime}_{\langle n,a\rangle}(e_{\langle n,a-1\rangle}(\nu_{a}))=\nu_{\langle n,a\rangle}$. Put $e_{\langle n,a\rangle}:=e^{\prime}_{\langle n,a\rangle}\circ e_{\langle n,a-1\rangle}$.
@@
This completes the definition of the strategy §$\S$ in the weak grouped Menger game played on $X$. It remains to notice that any play in this game in which Alice uses §$\S$ gives a sequence of objects we require in our lemma, such that conditions $(1)$$(4)$ are satisfied, and if Alice loses (and such a play exists because §$\S$ cannot be winning), then also $(5)$ is satisfied. ∎

### v3

@@ 4. Totally imperfect Menger sets and Sacks forcing
For the proof of Theorem 4.1, we need the following auxiliary result, whose proof is rather standard (see, e.g., [6, Lemma 5.10] for a similar argument) and is left to the reader.
@@
Take $p$ and $F_{n},k_{n},l_{n},y_{\sigma_{n}}$ from Lemma 3.7, applied to $p_{0}$ and $\tau$. Note also that Lemma 3.9(3) ensures that $S=\mathrm{supp}(p)$ and the sequences ${\langle\,k_{n}:n\in\omega\,\rangle},{\langle\,F_{n}:n\in\omega\,\rangle},{\langle\,\Sigma_{n}:n\in\omega\,\rangle}$ satisfy $(e_{f})$ from the first paragraph of Section 2. Let $K$ and $\sigma]$ for $\sigma\in\bigcup_{n\in\omega}\Sigma_{n}$ be defined in the same way as in the first paragraph of Section 2.
@@
[…]
@@ 5. A modification of the Menger game and consonant spaces
Fix a natural number $n>0$ and assume that natural numbers $l_{0},\dotsc,l_{n-1}>0$ and n1n-1 rounds of the game have been defined.
@@
Let $\sigma$ be a strategy for Alice in the grouped Menger game played on $X:={2^{\omega}}\setminus Y$. By Remark 5.1, we may assume that each family given by Alice according to the strategy $\sigma$ is countable and increasing and Bob chooses one set from the families given by Alice. We shall define a strategy $\S$ for Alice in the game $\mathsf{G}_{1}(\mathcal{K},\mathcal{O})$ played on $X$ such that each play in $\mathsf{G}_{1}(\mathcal{K},\mathcal{O})$ played according to $\S$ and lost by Alice, gives rise to a play in the grouped Menger game on $X$ in which Alice uses $\sigma$ and loses.
Suppose that $\sigma$ instructs Alice to start round 0 by selecting a natural number $l_{0}>0$. Let $L_{0}:=0$ and $L_{1}:=l_{0}$. Then $\S$ instructs Alice to play the family of all sets
@@
is a play, where Alice uses the strategy $\sigma$. Note that the family of all the intersections as above is indeed an open $k$-cover of $X$. Suppose that Bob replies in $\mathsf{G}_{1}(\mathcal{K},\mathcal{O})$ by selecting $\bigcap_{i\in[L_{0},L_{1})}U_{i}$ for some sequence $(\mathcal{U}_{0},U_{0},\dotsc,\mathcal{U}_{l_{0}-1},U_{l_{0}-1})$ as above. The strategy $\sigma$ instructs Alice to proceed in round $1$ by selecting a natural number $l_{1}>0$. Let $L_{2}:=l_{0}+l_{1}$. Then $\S$ instructs Alice to play the family of all sets
@@
In general, let $\sigma$ instruct Alice to start round $n$ by selecting a natural number $l_{n}>0$. Let $L_{n+1}:=l_{0}+l_{1}+\dotsb+l_{n}$. Then the next move of Alice in $\mathsf{G}_{1}(\mathcal{K},\mathcal{O})$ according to $\S$ is, by the definition, the family of all sets
@@
Since the strategy $\S$ is not winning, there is a play in $\mathsf{G}_{1}(\mathcal{K},\mathcal{O})$ in which Alice uses $\S$ and loses, which gives rise to an infinite play
@@ 5.1. Menger game versus the grouped Menger game
Combining [10, Theorem 1.1] and [7, Theorem 10], we conclude that $\mathfrak{d}=\mathfrak{c}$ implies the existence of a Menger ultrafilter.
@@
Let us note that there are models of ZFC without Menger ultrafilters. Indeed, every Menger ultrafilter is a $P$-point, see, e.g., [5, Observation 3.4]. According to a result of Shelah published in [28] (see also [8]), consistently there are no $P$-points.
Next, we introduce a variation of the grouped Menger game which is “harder” for Alice since her choices are more restricted, i.e., are not arbitrary open covers of the space in question. Let $K\subseteq{2^{\omega}}$ be a perfect set. For $y\in K$, let $\mathcal{U}_{y}$ be a countable increasing family of clopen sets in $K$ such that $\bigcup\mathcal{U}_{y}=K\setminus\{y\}$. Let $X\subseteq K$ be a set such that $K\setminus X$ is dense in $K$. The weak grouped Menger game played on $X$ in $K$ ($\operatorname{wgM}(K,X)$ in short) is played as follows: In round 0 Alice selects a natural number $l_{0}>0$, and Bob selects a closed nowhere dense subset $K_{0}$ of $K$. Let $L_{0}:=0$ and $L_{1}:=l_{0}$. Then the players play the usual Menger game $l_{0}$ subrounds, with the following restrictions. In each subround $i\in[L_{0},L_{1})$, Alice chooses $y_{i}\in K\setminus(X\cup K_{0})$ and plays the family $\mathcal{U}_{i}:=\mathcal{U}_{y_{i}}$. Then Bob replies by choosing a set $U_{i}\in\mathcal{U}_{i}$ with $K_{0}\subseteq U_{i}$.
@@
$\textrm{B}\uparrow\operatorname{wgM}(K,X\cap K)$$\textrm{A}\not\uparrow\operatorname{wgM}(K,X\cap K)$$X\text{ is perfectly meager}$${2^{\omega}}\setminus X\text{ is consonant}$$X\text{ is Hurewicz}$$X\text{ is Rothberger}$$X\text{ is Menger}$}
@@
We shall describe a strategy $\S$ for Alice in the weak grouped Menger game played on $X$ in $K$ such that each play lost Alice gives rise to the objects whose existence we need to establish.
Round 0. Let $C_{0}:=\Sigma_{0}$ and $i_{0}:=0$. Alice declares that the 0-th group will have length $l_{0}:=|C_{0}|$. Let $\{\nu_{j}:j<l_{0}\}\subseteq\Sigma_{0}$ be an enumeration of $C_{0}$. Suppose that Bob plays a closed nowhere dense $K_{0}\subseteq K$.
Subround $\langle 0,0\rangle$. By density of $K\setminus X$, Alice picks $y_{\langle 0,0\rangle}\in[\nu_{0}]\setminus(X\cup K_{0})$ and she plays the family $\mathcal{U}_{\langle 0,0\rangle}:=\mathcal{U}_{y_{\langle 0,0\rangle}}$. Suppose that Bob replies by choosing $U_{\langle 0,0\rangle}\in\mathcal{U}_{\langle 0,0\rangle}$ with $K_{0}\subseteq U_{\langle 0,0\rangle}$. Take $j_{\langle 0,0\rangle}>i_{0}$ and $\nu_{\langle 0,0\rangle}\in\Sigma_{j_{\langle 0,0\rangle}}$ with $\nu_{\langle 0,0\rangle}\succ\nu_{0}$ such that $y_{\langle 0,0\rangle}\in[\nu_{\langle 0,0\rangle}]$ and $\nu_{\langle 0,0\rangle}]\cap U_{\langle 0,0\rangle}=\emptyset$. Let $e_{\langle 0,0\rangle}\colon C_{0}\to\Sigma_{j_{\langle 0,0\rangle}}$ be a coherent map such that $e_{\langle 0,0\rangle}(\nu_{0})=\nu_{\langle 0,0\rangle}$. Fix a natural number $a$ with $0<a<l_{0}$ and assume that the players have already defined the following sequences:
@@
  • $\langle\,\nu_{\langle 0,b\rangle}:0\leq b<a\,\rangle$} of maps $\nu_{\langle 0,b\rangle}\in\Sigma_{j_{\langle 0,b\rangle}}$ such that $y_{\langle 0,b\rangle}\in[\nu_{\langle 0,b\rangle}]$, $\nu_{b}\prec\nu_{\langle 0,b\rangle}$ and $\nu_{\langle 0,b\rangle}]\cap U_{\langle 0,b\rangle}=\emptyset$,
@@
Subround $\langle 0,a\rangle$: Alice picks $y_{\langle 0,a\rangle}\in[e_{\langle 0,a-1\rangle}(\nu_{a})]\setminus(X\cup K_{0})$ and she plays the family $\mathcal{U}_{\langle 0,a\rangle}:=\mathcal{U}_{y_{\langle 0,a\rangle}}$. Suppose that Bob replies by choosing $U_{\langle 0,a\rangle}\in\mathcal{U}_{\langle 0,a\rangle}$ with $K_{0}\subseteq U_{\langle 0,a\rangle}$. Take $j_{\langle 0,a\rangle}>j_{\langle 0,a-1\rangle}$ and $\nu_{\langle 0,a\rangle}\in\Sigma_{j_{\langle 0,a\rangle}}$ with $\nu_{\langle 0,a\rangle}\succ e_{\langle 0,a-1\rangle}(\nu_{a})$ such that $y_{\langle 0,a\rangle}\in[\nu_{\langle 0,a\rangle}]$ and $\nu_{\langle 0,a\rangle}]\cap U_{\langle 0,a\rangle}=\emptyset$. Let $e^{\prime}_{\langle 0,a\rangle}\colon e_{\langle 0,a-1\rangle}[C_{0}]\to\Sigma_{j_{\langle 0,a\rangle}}$ be a coherent map such that $e^{\prime}_{\langle 0,a\rangle}(e_{\langle 0,a-1\rangle}(\nu_{a}))=\nu_{\langle 0,a\rangle}$. Put $e_{\langle 0,a\rangle}:=e^{\prime}_{\langle 0,a\rangle}\circ e_{\langle 0,a-1\rangle}$.
@@
satisfy all relevant instances of (1)(1)$(4)$.
@@
Subround $\langle n,0\rangle$: Alice picks $y_{\langle n,0\rangle}\in[\nu_{0}]\setminus(X\cup K_{n})$ and she plays the family $\mathcal{U}_{\langle n,0\rangle}:=\mathcal{U}_{y_{\langle n,0\rangle}}$. Suppose that Bob replies by choosing $U_{\langle n,0\rangle}\in\mathcal{U}_{\langle n,0\rangle}$ with $K_{n}\subseteq U_{\langle n,0\rangle}$. Take $j_{\langle n,0\rangle}>i_{n}$ and $\nu_{\langle n,0\rangle}\in\Sigma_{j_{\langle n,0\rangle}}$ with $\nu_{\langle n,0\rangle}\succ\nu_{0}$ such that $y_{\langle n,0\rangle}\in[\nu_{\langle n,0\rangle}]$ and $\nu_{\langle n,0\rangle}]\cap U_{\langle n,0\rangle}=\emptyset$. Let $e_{\langle n,0\rangle}\colon C_{n}\to\Sigma_{j_{\langle n,0\rangle}}$ be a coherent map such that $e_{\langle n,0\rangle}(\nu_{0})=\nu_{\langle n,0\rangle}$.
@@
  • $\langle\,\nu_{\langle n,b\rangle}:0\leq b<a\,\rangle$} of maps $\nu_{\langle n,b\rangle}\in\Sigma_{j_{\langle n,b\rangle}}$ such that $y_{\langle n,b\rangle}\in[\nu_{\langle n,b\rangle}]$, $\nu_{b}\prec\nu_{\langle n,b\rangle}$ and $\nu_{\langle n,b\rangle}]\cap U_{\langle n,b\rangle}=\emptyset$;
@@
Subround $\langle n,a\rangle$: Alice picks $y_{\langle n,a\rangle}\in[e_{\langle n,a-1\rangle}(\nu_{a})]\setminus(X\cup K_{n})$ and she plays the family $\mathcal{U}_{\langle n,a\rangle}:=\mathcal{U}_{y_{\langle n,a\rangle}}$. Suppose that Bob replies by choosing $U_{\langle n,a\rangle}\in\mathcal{U}_{\langle n,a\rangle}$ such that $K_{n}\subseteq U_{\langle n,a\rangle}$. Take $j_{\langle n,a\rangle}>j_{\langle n,a-1\rangle}$ and $\nu_{\langle n,a\rangle}\in\Sigma_{j_{\langle n,a\rangle}}$ with $\nu_{\langle n,a\rangle}\succ e_{\langle n,a-1\rangle}(\nu_{a})$ such that $y_{\langle n,a\rangle}\in[\nu_{\langle n,a\rangle}]$ and $\nu_{\langle n,a\rangle}]\cap U_{\langle n,a\rangle}=\emptyset$. Let $e^{\prime}_{\langle n,a\rangle}\colon e_{\langle n,a-1\rangle}[C_{n}]\to\Sigma_{j_{\langle n,a\rangle}}$ be a coherent map such that $e^{\prime}_{\langle n,a\rangle}(e_{\langle n,a-1\rangle}(\nu_{a}))=\nu_{\langle n,a\rangle}$. Put $e_{\langle n,a\rangle}:=e^{\prime}_{\langle n,a\rangle}\circ e_{\langle n,a-1\rangle}$.
@@
This completes the definition of the strategy $\S$ in the weak grouped Menger game played on $X$. It remains to notice that any play in this game in which Alice uses $\S$ gives a sequence of objects we require in our lemma, such that conditions (1)(1)$(4)$ are satisfied, and if Alice loses (and such a play exists because $\S$ cannot be winning), then also $(5)$ is satisfied. ∎